# Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES)

### v2

@@ IV-E Annotation using HRI CUES
After establishing HRI CUES, all annotators independently rated the remaining 25 videos described in Section III-B. The same methodology was employed as in Section IV-B, with the only difference being the enjoyment scale, as HRI CUES was used instead of the initial scale. That is, the annotators rated 590 turns (174 min) using ELAN with both side and frontal view of the interaction (Fig. 2), viewing each turn only once, in addition to providing an overall enjoyment score per interaction. They also provided an explanation for their overall ratings and the challenges they faced during the annotation, as in Section IV-B. The annotation was conducted over 8 days. The results are reported in the next section.
@@ V Results
Following the discussions involved in the annotator alignment that redefined the user enjoyment scale and methodology, annotators rated the remaining 25 videos of robot interactions from our participatory design workshop individually.
@@ V-A Distribution of User Enjoyment
Fig. 3 shows how each annotator rated the interaction exchanges (turns), indicating that the interactions mainly were (%) regarded as neutral in enjoyment, with rare occurrences of very low (%) and very high (%) enjoyment, showing a near Gaussian distribution of user enjoyment for each annotator. Fig. 8 (in the Appendix) shows the rating distributions of annotators per participant, which display a similar Gaussian distribution of perceived enjoyment by the annotators, with some participants (e.g., P25, P27) perceived to have a more enjoyable interaction than others (e.g., P3).
@@ V-B Rater Reliability
Similarly to how our data was coded, we present rater reliability for each conversation exchange and the overall enjoyment score provided by each annotator for the interactions.
@@ V-B1 Per Conversation Turn
The resulting annotations per conversation turn of 25 videos are shown in Fig. 9 in the Appendix. Treating each conversation turn as a repeated measures factor in the reliability analysis:
@@
These results confirmed our initial conclusion. In addition, removing A2 or A3 separately decreased ICC. Subsequent discussions with the annotators further confirmed a divergence in the ratings provided by A1 (Section VI-C2).
@@ V-B2 Overall Enjoyment Score
V-C Correlations with Self-Reported User Perceptions
Users’ subjective ratings of enjoyment during the interactions were obtained from the questionnaire in the participatory design workshops after their interaction with the robot, in terms of user satisfaction, fun, interestingness, and strangeness of the conversation, as described in Section III-B, which are presented in Fig. 2 (in the Appendix) along with the annotator overall enjoyment ratings per interaction. The items had high correlation (Cronbach’s alpha = ), with the removal of each item reducing the correlation in the construct. These self-reported scores and the average of these scores were compared against annotators’ overall enjoyment scores to evaluate how well the annotators could perceive their enjoyment. Spearman correlation was used across four Likert scale items and the average of these scores, with 95% confidence interval (ranging from to ). The results were as follows:
@@ VI-A Enhanced Reliability with Averaged Annotator Ratings
The superior reliability of ICC(2,k) highlights the inherent variability in subjective experiences and perceptions of enjoyment, suggesting that averaging across multiple annotators can effectively mitigate individual biases and variations in judgment, leading to a more reliable representation of true user enjoyment. This finding is critical, as it emphasizes the importance of incorporating multiple perspectives to achieve a more accurate and consistent evaluation of user enjoyment in conversational interactions with robots. Consequently, the distinction between ICC(2) and ICC(2,k) results not only validates the robustness of our user enjoyment scale but also illustrates the methodological importance of employing multiple annotators for capturing the complex and subjective nature of enjoyment in human-robot conversations.
VI-B Correlation with Self-rated Enjoyment Scale
@@
To further validate our user enjoyment scale, we analyzed correlations between the annotators’ overall enjoyment scores and the participants’ subjective enjoyment ratings. Results revealed a statistically significant, moderate positive correlation between overall enjoyment scores and the (reversed) ‘felt strange’ item, indicating that higher enjoyment scores were associated with decreased feelings of awkwardness during the interaction. This appears to greatly align with our annotator’s discussions and the resulting enjoyment scale that classifies the presence of signs of discomfort as the scale’s lowest level.
@@
However, the correlations between annotator scores and other user-reported measures — such as satisfaction, average enjoyment, fun in talking, and interest in conversation — although trending towards moderate positive correlations, did not reach statistical significance. These findings suggest a nuanced relationship between observed enjoyment and user self-reported experiences. While annotators can detect general levels of comfort and ease within interactions, capturing comprehensive internal subjective enjoyment may require additional data, such as initial expectations towards the robot, personality traits, as well as physiological responses, not accessible through direct observation. This discrepancy underscores the complexity of correlating observed behavior with subjective internal states and highlights the challenge of fully capturing user enjoyment in HRI. Despite these considerations, having an additional tool to evaluate user enjoyment in conversations is highly valuable for HRI research to assess and develop robots that are enjoyable to interact with, within all parts of the conversation, especially in daily life. This approach aligns with how humans naturally adapt our conversations in real-time based on the external cues we observe in others. The measure presented in this paper offers a valuable tool for collecting observational data to train autonomous enjoyment detection systems that can be used on robots or other agents.
VI-C Final Alignment Post Annotation
After analyzing the results, the annotators were asked to have another set of discussions by presenting them with Fig. 9 and 10 in the Appendix (without reliability or correlation scores), which lasted four hours. They were asked to pinpoint instances of rating divergence and concordance within their assessments. Particular emphasis was placed on identifying turns where annotators disagreed or agreed most fervently. Subsequently, the annotators were prompted to watch the corresponding video segments to explore the reasons behind their ratings and the rationale for their agreement or disagreement. The primary objective was to gain insight into potential major concordance (Section VI-C1) and divergence (Section VI-C2) in the way dialogue exchanges were annotated and elucidate the underlying reasons for these variations.
Following the turn-by-turn analysis, annotators were asked to identify the two most significant discrepancies (highest and lowest) between their ratings and self-reported participant perceptions (Section VI-C3). They were asked to engage in discussions exploring potential reasons underlying the disparities and similarities between their perceptions and those of the users from multiple aspects based on their expertise.
VI-C1 Concordance Between Annotators
In numerous instances across the videos, a concordant agreement was observed among the annotators. As an illustrative example, in the interaction of the ninth participant (P9, turn 13 in Fig. 4), the conversation exchange exhibited a seamless progression, and the participant’s enjoyment level was distinctly conveyed through expansive bodily gestures. Notably, all annotators unanimously assigned a rating of five for the exchange since the participant threw themselves backward in the chair laughing. In the same video, the annotators all assigned a rating of two for the exchanges where the participant’s response was marked by sighing and a demeanor suggestive of resignation. This reaction occurred as a response to an unnecessary repetition initiated by the robot, specifically at turn 23.
In another example (P10, Fig. 4), the annotators assigned a rating of 1 to turn 16 to indicate that the participant openly expressed their negative thoughts due to the robot not making eye contact with the participant. During this interaction, the participant also attempted to establish contact with the experimenter. Subsequently, the participant made an effort to politely maintain a dialogue with the robot according to social norms, a behavior that garnered consensus among the annotators as being representative of a rating of 3.
In conclusion, the annotators agreed when the user enjoyment scale aligned clearly with participant behavior. However, in most cases, the interaction between the robot and the participant did not correspond as clearly or unambiguously to the user enjoyment scale. This is likely due to the complex and situational nature of the cues in the interaction, making it challenging to develop comprehensive yet precise guidelines for annotation. Instead, the general scale needs to be interpreted by the annotators for the particular use case to find anchor points that are appropriate for the specific context.
VI-C2 Divergence Between Annotators
Throughout the analysis of the 25 videos, there were instances where annotators differed substantially in their assessments. For instance, for P1 (Fig. 5), at turn 4, A1 assigned a rating of 5, while A2 scored it as 2, and A3 as 3. The participant’s laughter posed a challenge as it was perceived both as a sign of high enjoyment (by A1) and, conversely, as an expression of frustration towards the situation or the robot (by A2 and A3), rather than amusement with the robot. Furthermore, at turn 15, A1 assigned a rating of 3, while A2 rated it 1, and A3 as 2. In this context, the participant remained entirely silent, awaiting the robot to initiate further interaction. Annotators interpreted this silence differently, seeing it as politeness, boredom, or discomfort.
For another participant (P25, Fig. 5), during turn 8, A1 assigned a rating of 5, whereas A2 rated it as 1, and A3 as 2. The participant asked the robot to make more eye contact with them, which could be interpreted as a period of heightened immersion and anthropomorphism or criticism. Following this, the annotators consistently exhibited discord in their assessments until turn 17, when they reached a consensus once more. For instance, at turn 11, A1 marked it as a 5, A2 as 3, and A3 as 1. In this case, the participant expressed reservations about sharing personal information with the robot due to unfamiliarity, yet did so while smiling and posing a question to the robot in a playful tone. This complexity in the interaction exchange presented challenges for the annotators, as it encompassed a multitude of actions. While the verbal content suggested discomfort, the presence of laughter, smiling, and playful tonality indicated enjoyment. Consequently, the annotators encountered mixed signals, and the resulting ratings depended on which aspect of the interaction they prioritized.
The notable inconsistencies between A1’s ratings and those of the others led to the inclusion of reliability results for the more consistently aligned group of annotators (A2 and A3).
VI-C3 Similarities and Discrepancies Between User and Annotator Perceptions
While annotators aligned well with a large proportion of the participants in their perceptions, there were substantial discrepancies for some of the participants. For instance, for P6 and P21 (see Fig. 6), it can be noted that the annotators and the participant interpreted user enjoyment in a similar way, with overall interaction scored as 5 and 4, matching that of the average of user reported values. The conversations went well, and the participants seemed to take a playful approach in the interaction, which was reflected in both the participants’ and annotators’ scores.
The conversation context might cause a discrepancy between the annotators’ assessment and users’ rating of the enjoyment. For instance, P3 (Fig. 7) talked about a controversial topic (UFOs) with the robot. The robot repeatedly questioned the participant (e.g., “Why do you think that?”, “Can you tell more about what gave you this insight?”) when they were affirmative about having observed the existence of UFOs. The repetitive questions could have been perceived as offensive or discomforting due to the nature of the topic and their stance towards it, despite it being a type of interaction failure (repetition of the same phrase) that occurred with other participants as well. The participant stated to the robot that they wanted to change the topic twice, and then turned to the experimenters to voice this desire (after 3.5 minutes), in addition to displaying cues of anxiety (e.g., playing with fingers, looking around at the camera and at the experimenters), which can confirm the belief from the annotators that the participant had a negative experience, who rated the interaction low in enjoyment. The participant managed to change the topic on their own to talk about the weather, and later about the university and robots for the rest of the conversation. However, the participant gave high scores (see Fig. 7) in all aspects of enjoyment. The participant might have rated the experience as more enjoyable and interesting than the annotators due to researcher bias, i.e., that the participant felt the need to please the researcher [64]. Given the controversial topic discussed, the positive ratings can be interpreted as a strategy to avoid judgment from the researchers, as the participant frequently gazed at the experimenter during several exchanges, while displaying signs of enjoyment (e.g., smiles, thumbs up, nods) even after the topic change. Another reason could be the novelty effect, since they might have been happy to talk with a robot regardless of the negative experience.
Another participant (P13, Fig. 7) experienced the interaction as less enjoyable than the annotators interpreted. This might be because the participant was experiencing a high number of technical and social failures from the robot while still displaying enjoyable signs. The participant used the robot for transactional requests rather than a casual conversation. The transactional nature of the conversation combined with failures might explain why the participant gave low enjoyment scores. The annotators gave a higher score because the participant seemed to forgive the failures, laugh them off, and continue the conversation smoothly. This can be seen as an important reminder that the annotators are not always assessing the same aspects as the participants in their self-assessment.
VI-D Adapting HRI CUES to Other Domains
@@
To replicate our methodology and adapt HRI CUES to another domain, the following framework should be applied:
@@ VI-E Challenges and Limitations
In this work, we introduced a novel scale for annotators to evaluate user enjoyment in conversational HRI. However, enjoyment is a subjective measure, and thus, is challenging to evaluate and agree on between annotators and correlate with users perceptions, given the multitude of aspects connected to it. For instance, the context and length of the exchange affected how the enjoyment was perceived. When segments were overly brief, the exchange did not always contain sufficient information for a fair assessment. More frequently, however, excessively long segments were complex to analyze. For example, as described in Section VI-C2, the annotators often interpreted the exchange differently due to long segments that contained several cues belonging to separate levels of the scale. In these cases, a different assessment approach may be necessary as the longer segments introduced an additional factor for the annotators to consider: which part or aspect of the exchange to emphasize in the assessment. Due to the complex nature of enjoyment as a concept, which highlights the necessity of pre-coding discussions among the annotators, aligning on the definition of enjoyment of relevance to the particular study is challenging. One solution is to focus on the change of behavior within the exchange. For instance, if the robot’s response improved the user’s demeanor towards the robot, the exchange should be rated towards the level that contains higher enjoyment cues in the scale, and conversely if it had a negative impact. Other alternatives could be to use an aggregated score, the rating that corresponds to the majority of the segment, the most/least extreme rating, or the rating that corresponds to the first or last part of the segment. The choice of method for these situations should be determined during the annotation alignment process, taking into account the specific use case and domain.
@@
The differences between the annotators’ overall rating on user enjoyment and user perceptions (as discussed in Section VI-C3) might be due to the participant’s expectation of the robot’s social and technical level, while the annotators only look at the interaction itself. In addition, since this is their first interaction with a robot for most of the participants (20 out of 25), the ‘novelty effect’ might have changed their perceptions more positively or negatively, given that the duration (7 minutes) is not long enough to overcome it [65]. However, the variability observed by the annotators in user enjoyment states throughout the interaction (e.g., Fig. 5) shows that conversation context may alter the novelty effect, providing a more complete picture of the enjoyment throughout the interaction than a self-reported score at the end of the interaction. In addition, the users’ responses to the questionnaire may differ from their actual attitudes towards the robot [66]. These support the importance of HRI CUES as an additional tool to evaluate user enjoyment, providing means for real-time estimation of enjoyment in conversational agents.
@@
Our scale is designed to serve as a tool for assessing perceived user enjoyment during interactions with robots, intended primarily for researchers in the field of HRI. However, given that open-domain dialogue may involve sensitive information that individuals may be hesitant to share with unfamiliar parties, it is imperative to obtain explicit consent from participants before they talk with the robot. In our studies, we ensured this consent prior to participants’ interactions with the robot, employing the use of the term ‘sentiment analysis’ in the consent form, and explained to all participants that their interactions would be analyzed in terms of their affective states (‘identification of feelings’) by both researchers and automated systems. While obtaining consent is essential for researchers utilizing our scale in future studies, it is crucial to recognize that this process may influence participants’ behavior and conversation topics, as they may be reluctant to share sensitive memories or be concerned about being judged by others. Consequently, this can result in a disconnection with the robot, posing challenges in achieving high levels of enjoyment during interactions. Nonetheless, this impact is likely to diminish over the course of the interaction or across multiple interactions, particularly in long-term settings. Additionally, researchers must exercise caution to uphold the privacy and confidentiality of participants when sharing data, ensuring that they remain unidentifiable in images and videos (as demonstrated in the examples provided for the scale), and removing any sensitive information. Moreover, researchers should remain mindful of their own biases and subjectivity, which may lead to variations in the interpretation of enjoyment compared to the participants’ experiences. User enjoyment is often context-specific, indicating that users’ behavioral and affective expressions are connected to specific socio-cultural contexts, including values, norms, and expectations of what is considered appropriate in certain situations [67]. These underscore the significance of employing multiple annotators for the scale that are familiar with the socio-cultural background of the target population, and using the scale to complement self-reported user perceptions to have a deeper understanding of the interaction quality. As human-robot interactions continue to evolve, ensuring a deep understanding of user enjoyment not only elevates the quality of these interactions but also paves the way for more empathetic and meaningful connections.

### v3

@@ IV-F Annotation using HRI CUES
After establishing HRI CUES, all annotators independently rated the remaining 25 videos described in Section III-D. The same methodology was employed as in Section IV-B, with the only difference being the enjoyment scale, as HRI CUES was used instead of the initial scale. That is, the annotators rated 590 turns (174 minutes) using ELAN with both side and frontal view of the interaction (Fig. 2), viewing each exchange only once, in addition to providing an overall enjoyment score per interaction. They also provided an explanation for their overall ratings and the challenges they faced during the annotation, as in Section IV-B. The annotation was conducted over 8 days. The results are reported in the next section.
@@ V Results
Following the discussions involved in the annotator alignment that redefined the user enjoyment scale and methodology, annotators rated the remaining evaluation data (25 videos) individually.
@@ V-A Distribution of User Enjoyment
Fig. 3 shows how each annotator rated the interaction exchanges (robot-participant turn pairs), indicating that the interactions mainly were (%) regarded as neutral in enjoyment, with rare occurrences of very low (%) and very high (%) enjoyment, showing a near Gaussian distribution of user enjoyment for each annotator. Fig. 4 (in the Appendix) shows the rating distributions of annotators per participant, which display a similar Gaussian distribution, with some participants (e.g., P25, P27) perceived to have a more enjoyable interaction than others (e.g., P3).
@@ V-B Rater Reliability
Similarly to how our data was coded, we present rater reliability for each conversation exchange and the overall enjoyment score provided by each annotator for the interactions. Detailed analysis of concordance and divergence between annotators is provided in the Appendix.
@@ V-B1 Per Conversation Turn
The resulting annotations per conversation turn of 25 videos are shown in Fig. 5 in the Appendix. Treating each exchange as a repeated measures factor in the reliability analysis:
@@
These results confirmed our initial conclusion. In addition, removing A2 or A3 separately decreased ICC. Subsequent discussions with the annotators further confirmed a divergence in the ratings provided by A1 (Section -B).
@@ V-B2 Overall Enjoyment Score
V-C Annotator Correlations with Self-Reported Enjoyment
Users’ subjective ratings of enjoyment during the interactions were obtained from the questionnaire in the second study after their interaction with the robot, in terms of user satisfaction, fun, interestingness, and strangeness of the conversation, as described in Section III-B2, which are presented in Fig. 8 (in the Appendix) along with the annotator overall enjoyment ratings per interaction. The items had high reliability (Cronbach’s alpha = ), with the removal of each item reducing the correlation in the construct. These self-reported scores and the average of these scores were compared against annotators’ overall enjoyment scores to evaluate how well the annotators could perceive their enjoyment. Spearman correlation was used across four Likert scale items and the average of these scores, with 95% confidence interval (ranging from to ). The results were as follows:
@@ V-C Annotator Correlations with Self-Reported Enjoyment
Detailed analysis of similarity and discrepancy between annotators and user perceptions is provided in the Appendix Section -C.
V-D Large Language Model Correlations with Self-Reported Enjoyment
HRI CUES was further examined for its potential in automatic enjoyment detection. To this end, the turn-by-turn enjoyment ratings generated by various LLM configurations (for more details see [20]) using the scale were averaged for each interaction and correlated with the same participants’ post-interaction self-reported ratings.
The results show a more notable alignment between LLM assessments and user perceptions. The GPT-4 Turbo model configured with all prompt features (including scale, instructions, examples, conversation history, and Chain-of-Thought reasoning) achieved the following statistically significant correlations with user self-reported ratings:
LLM Average vs. User Average: Statistically significant () and strong positive correlation ().
LLM Average vs. User Satisfaction: Statistically significant () and strong positive correlation ().
LLM Average vs. User Fun Talking: Statistically significant () and moderate positive correlation ().
LLM Average vs. User Conversation Interesting: Not statistically significant () moderate positive correlation ().
LLM Average vs. User Felt Strange (Reversed): Statistically significant () and moderate positive correlation () .
Other LLM configurations, including multimodal models like Gemini Pro 1.5, also showed significant moderate correlations, particularly with satisfaction and strangeness.
LLMs have also shown the ability to predict turn-by-turn user enjoyment with HRI CUES even without access to the user’s response to a robot’s utterance [21]. The turn-by-turn exchange level predictive capability performs comparably to enjoyment detection (with user response) from the human experts annotations reported here. On a correlation level, even when removing user responses, GPT‑4o still achieved significant correlation values of () with satisfaction and () with strangeness (reversed).
@@ VI-A Enhanced Reliability with Averaged Annotator Ratings
The superior reliability of ICC(2,k) highlights the inherent variability in subjective experiences and perceptions of enjoyment, suggesting that averaging across multiple annotators can effectively mitigate individual biases and variations in judgment. Divergence of annotator perceptions based on the context and cues of the exchange (detailed in Section -B in Appendix) aligns with challenges detected for user satisfaction annotations [12], suggesting that averaging across annotators can lead to a more reliable representation of true user enjoyment. This finding is critical, as it emphasizes the importance of incorporating multiple perspectives to achieve a more accurate and consistent evaluation of user enjoyment in conversational interactions with robots. Consequently, the distinction between ICC(2) and ICC(2,k) results not only supports the robustness of our user enjoyment scale but also illustrates the methodological importance of employing multiple annotators for capturing the complex and subjective nature of enjoyment in human-robot conversations.
VI-B Correlation with User Perceptions
@@
To further evaluate our user enjoyment scale, we analyzed correlations between the annotators’ overall enjoyment scores and the participants’ subjective enjoyment ratings. Results revealed a statistically significant, moderate positive correlation between overall enjoyment scores and the (reversed) ‘felt strange’ item, indicating that higher enjoyment scores were associated with decreased feelings of awkwardness during the interaction. This appears to greatly align with our annotator’s discussions and the resulting enjoyment scale that classifies the presence of signs of discomfort as the scale’s lowest level.
@@
However, the correlations between annotator scores and other user-reported measures — such as satisfaction, average enjoyment, fun in talking, and interest in conversation — although trending towards moderate positive correlations, did not reach statistical significance. These findings suggest a nuanced relationship between observed and user experiences (detailed in Section -C in Appendix). While annotators can detect general levels of comfort and ease within interactions, capturing comprehensive internal subjective enjoyment may require additional data, such as initial expectations towards the robot, personality traits, as well as physiological responses, not accessible through direct observation. This discrepancy underscores the complexity of correlating observed behavior with subjective internal states and highlights the challenge of fully capturing user enjoyment in HRI.
However, complementary evidence comes from our LLM analysis. GPT-4 Turbo, guided by the full HRI CUES prompt, achieved strong and statistically significant correlations with most self-ratings, including for the user-average score, and even the GPT-4o variant—without access to the user’s replies—retained significant correlations. Multimodal models such as Gemini Pro 1.5 showed comparable, moderately strong alignments. These results indicate that evaluating enjoyment with HRI CUES using automated systems, such as foundation models, can reveal subjective enjoyment cues that human observers may overlook, likely due to the models’ ability to integrate subtle lexical, prosodic, and contextual patterns throughout the dialogue. This finding aligns with prior literature on automatic assessment outperforming annotators in correlations with user perceptions [13]. The convergence between self-ratings and LLM predictions points to a path toward real-time, automated enjoyment monitoring during open-domain dialogue in HRI.
Our approach aligns with how humans naturally adapt our conversations in real-time based on the external cues we observe in others. The measure presented in this paper offers a valuable tool for collecting observational data to train autonomous enjoyment detection systems that can be used on robots or other agents.
VI-C Adapting HRI CUES to Other Domains
@@
To replicate our methodology and adapt HRI CUES to another domain (assuming that a clear understanding of the intended users, domain, and context is already established), the following framework should be applied:
@@ VI-D Challenges and Limitations
In this work, we introduced a novel scale for annotators to evaluate user enjoyment in conversational HRI. However, enjoyment is a subjective measure, and thus, is challenging to evaluate and agree on between annotators and correlate with users perceptions, given the multitude of aspects connected to it. For instance, the context and length of the exchange affected how the enjoyment was perceived. When segments were overly brief, the exchange did not always contain sufficient information for a fair assessment. More frequently, however, excessively long segments were complex to analyze. For example, the annotators often interpreted the exchange differently due to long segments that contained several cues belonging to separate levels of the scale (described in Section -B). In these cases, a different assessment approach may be necessary as the longer segments introduced an additional factor for the annotators to consider: which part or aspect of the exchange to emphasize in the assessment. Due to the complex nature of enjoyment as a concept, which highlights the necessity of pre-coding discussions among the annotators, aligning on the definition of enjoyment of relevance to the particular study is challenging. One solution is to focus on the change of behavior within the exchange. For instance, if the robot’s response improved the user’s demeanor towards the robot, the exchange should be rated towards the level that contains higher enjoyment cues in the scale, and conversely if it had a negative impact. Other alternatives could be to use an aggregated score, the rating that corresponds to the majority of the segment, the most/least extreme rating, or the rating that corresponds to the first or last part of the segment. The choice of method for these situations should be determined during the annotation alignment process, taking into account the specific use case and domain.
@@
The differences between the annotators’ overall rating on user enjoyment and user perceptions (detailed in Section -C in Appendix) might be due to the participant’s expectation of the robot’s social and technical level, while the annotators only look at the interaction itself. In addition, since this is their first interaction with a robot for most of the participants (20 out of 25), the ‘novelty effect’ might have changed their perceptions more positively or negatively, given that the duration (7 minutes) is not long enough to overcome it [77]. However, the variability observed by the annotators in user enjoyment states throughout the interaction (e.g., Fig. 7 in Appendix) shows that conversation context may alter the novelty effect, providing a more complete picture of the enjoyment throughout the interaction than a self-reported score at the end of the interaction. In addition, the users’ responses to the questionnaire may differ from their actual attitudes towards the robot [78]. These support the importance of HRI CUES as an additional tool to evaluate user enjoyment, providing means for real-time estimation of enjoyment in conversational agents.
A limitation worth noting was that the scale was developed and evaluated on open-domain conversations with a companion robot with older adults from Sweden. However, the scale development was initially based on enjoyment in marriage conversations between couples [18], in addition with various multimodal cues derived from the HRI and HCI literature in affective computing, user engagement, user enjoyment, conversation, and turn-taking analysis. In other words, HRI CUES levels or signs of enjoyment were not specific to older adults or the context of the interaction. Thus, the scale is generalizable to other contexts in which a user interacts with a social robot. In addition, we offered guidelines to adapt the scale to other application domains with robots or agents (Section VI-C). Nonetheless, future work is necessary to validate the scale in other domains [79], with a larger sample [80], participants from other cultures and socio-economic backgrounds [63, 79], a broader age range [81] and varying levels of familiarity with technology and robots [82], and in long-term deployments to mitigate the novelty effect [83], as users’ expectations and enjoyment of robot interactions may vary depending on individual differences, contextual factors, and the duration of interaction.
@@
Our scale is designed to serve as a tool for assessing perceived user enjoyment during interactions with robots, intended for researchers in the field of HRI and automatic systems for real-time or offline evaluation. However, given that open-domain dialogue may involve sensitive information that individuals may be hesitant to share with unfamiliar parties or automated systems, it is imperative to obtain explicit consent from participants before they talk with the robot. In our studies, we ensured this consent prior to participants’ interactions with the robot, employing the use of the term ‘sentiment analysis’ in the consent form, and explained to all participants that their interactions would be analyzed in terms of their affective states (‘identification of feelings’) by both researchers and automated systems. While obtaining consent is essential for researchers utilizing our scale in future studies, it is crucial to recognize that this process may influence participants’ behavior and conversation topics, as they may be reluctant to share sensitive memories or be concerned about being judged by others. Consequently, this can result in a disconnection with the robot, posing challenges in achieving high levels of enjoyment during interactions. Nonetheless, this impact is likely to diminish over the course of the interaction or across multiple interactions, particularly in long-term settings. Additionally, researchers must exercise caution to uphold the privacy and confidentiality of participants when sharing data, ensuring that they remain unidentifiable in images and videos (as demonstrated in the examples provided for the scale), and removing any sensitive information. Moreover, researchers should remain mindful of their own biases and subjectivity, which may lead to variations in the interpretation of enjoyment compared to the participants’ experiences. User enjoyment is often context-specific, indicating that users’ behavioral and affective expressions are connected to specific socio-cultural contexts, including values, norms, and expectations of what is considered appropriate in certain situations [84]. These underscore the significance of employing multiple annotators for the scale that are familiar with the socio-cultural background of the target population, and using the scale to complement self-reported user perceptions to have a deeper understanding of the interaction quality. As human-robot interactions continue to evolve, ensuring a deep understanding of user enjoyment not only elevates the quality of these interactions but also paves the way for more empathetic and meaningful connections.